A man invested $2000, part at 8% and part at 6%. He earned $144
from his investments. How much did he invest at 8%

Answer by mananth(16946)

Investment I 8.00% per annum
Investment II 6.00% per annum

x+y=2000 ------------------------1
8.00% x + 6.00% y = $144.00
Multiply by 100
8x+6y= $14,400.00 --------2

Multiply (1) by -8
we get

-8x-8y= -16000.00

Add this to (2)

-2y=-$1,600.00

divide by -2

y=$800.00 investment at 6.00%
Balance $1,200.00 investment at 8.00%
m.ananth@hotmail.com